# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ IV.2 Paramagnet to spin glass transition scaling for MAX-3-XORSAT
For $\kappa_{c}=1.29$, this function is well fit by $\Omega_{0}=a\sqrt{N}2^{-bN}$, where $b\simeq 0.14$, in decent agreement with the result for a mean-field $p=3$-spin ferromagnet derived in Jörg et al. (2010). We note also the hardness equivalence between dense and random sparse graphs drawn in Basso et al. (2022), though we emphasize that the results we derive here are not limited to completely random graphs. If we are in the diabatic regime for $t_{f}\ll 1/\Omega_{0}\left(N\right)$, the probability of finding the ground state after a sweep is $P_{\mathrm{GS}}\left(t_{f}\right)\simeq\Omega_{0}^{2}t_{f}/W$, where $W\sim O\left(N\right)$ is the energy range swept over in the evolution; for $t_{f}\propto N$ we thus have $P_{\mathrm{GS}}\left(t_{f},N\right)\simeq N2^{-2bN}$. Our numerical simulations for PPSPs, shown below, are in good agreement with this result. Note that this assumes a single avoided crossing and does not consider transverse field chaos, e.g. crossings late in evolution between states close in energy; such transitions can in principle be substantially more difficult Schiffer et al. (2023), but are not relevant to approximation hardness in this problem as argued earlier.
IV.3 Scaling of inter-valley tunneling in a $p=3$ quantum spin glass
This calculation is less easy to generalize to a spectrally folded Hamiltonian, however, where we have exponentially many competing ground states, clustered in a thin hyperspherical shell around the true minimum, at least in approximation-hard problems. So instead we will consider tunneling between two $p$-spin wells in an $N$ spin system, spaced $N/2$ flips apart, and show that the average per-state tunneling rate in the TMA formulation of spectrally folded optimization should have near-identical scaling. Our total $N$-spin Hamiltonian, consequently, is
$-\kappa\sum_{j}X_{j},$ (16)
[…]
@@ IV.3 Scaling of inter-valley tunneling in a $p=3$ quantum spin glass
  • We divide the system between primary spins, which flip between the classical minima, and secondary spins, which do not. We then compute the dressed states $\left|0_{D}\right>$ and $\left|1_{D}\right>$ that comprise all the primary flip sequences up to order $M/2$ away from each minimum. It is at this order that the two states have nonzero overlap.
@@
$\left|0_{D}\right>\equiv\left|0\right>+\sum_{j}\frac{\kappa}{u_{1,0}}X_{j}\left|0\right>+\sum_{j,k\left(j\neq k\right)}\frac{2\kappa^{2}}{u_{1,0}u_{2,0}}X_{j}X_{k}\left|0\right>$
$+\sum_{j,k,l\left(j\neq k\neq l\right)}\frac{3!\kappa^{3}}{u_{1,0}u_{2,0}u_{3,0}}X_{j}X_{k}X_{l}\left|0\right>+...$ (19)
The expression for $\left|1_{D}\right>$ is identical. We stop our expansion at order $M/2$, which is the lowest nontrivial order needed to connect the states. For simplicity we assume $M$ is even though the argument is easy to generalize to odd $M$ as well. Note that this state is not normalized, and in fact the norm of the state written above is exponentially large, so we will need to incorporate normalization corrections into the definition of the states. Thanks to the dressing of the states, we obtain a primary-spin energy splitting
$\frac{1}{2}\left(\left<0_{D}\right|+\left<1_{D}\right|\right)H_{P}\left(\left|0_{D}\right>+\left|1_{D}\right>\right)=2\Omega_{0}^{(p)},$
@@
$\mathcal{N}^{(p)}$ $=$ $\left(1+\sum_{k=1}^{M/2}\binom{M}{k}\left(\kappa^{k}k!\prod_{j=1}^{k}\frac{1}{u_{j,0}}\right)^{2}\right)^{-1}.$ (21)
@@
We now turn to the secondary spins. To introduce secondary spin corrections, consider a single secondary spin $j$ out of the $\left(N-M\right)\sim N/2$ total, whose bit value is the same in both classical minima. Since the same transverse field is acting on it as all other spins, when we consider the sum of all processes that connect the two minima, we can now divide them between those where $M/2$ primary spins flip from each minima to meet in the middle, with secondary spin $j$ unchanged, and a new set of processes where spin $j$ flips starting from each minimum and the two wavefunctions overlap at the set of states where $M/2$ primary spins have flipped along with $j$. The first set of processes is what was considered in Eq. IV.3; the second is new, and we want to calculate its matrix element. It is most useful to express these matrix elements as a ratio of the new term to the original, primary-spin-only process, since both decay exponentially in $M$. Let the primary spin perturbative matrix element to reach $M/2$ flips be $\xi_{M/2}$, so that
@@
Taking all of these effects into account yields a highly accurate prediction of the minimum gap scaling for a wide range of values for $p$ and $\kappa$, with only an $O\left(1\right)$ discrepancy in the prefactor and few percent discrepancies in the scaling exponent (empirically, Eq. IV.3 tends to slightly overestimate the decay compared to the exponent extracted from numerical diagonalization). We refer to the Appendix A for more details.
@@
IV.4 Achievable approximation ratio with spectrally folded trial minimum annealing

### v3

@@ IV.2.1 First stage scaling
For $\kappa_{c}=1.29$, this function is well fit by $\Omega_{0}=a\sqrt{N}2^{-bN}$, where $b\simeq 0.14$, in decent agreement with the result for a mean-field $p=3$-spin ferromagnet derived in Jörg et al. (2010). And despite being a very simple calculation, it is a quantitatively good predictor of the scaling for TAQC in finding ground states of PPSPs, as shown numerically later on in section V.3.
We emphasize that this prediction is only an approximate average, of course. And for a filtering choice of $x\geq 2$, a more sophisticated method is needed to compute $\tilde{u}\left(\kappa_{c},x,w,m\right)$, as we can no longer use a simple perturbation expansion about $G$ owing to the vanishing local gap. But it illustrates something important, and non-perturbative: the scaling of $\Omega_{0}$ is set by the amplitudes in $\left|G_{D}\right>$ of exponentially many configurations extensive distances away from $G$.
These amplitudes are functions of both $\kappa_{c}$–which we have already argued should be stable across most appropriately normalized instances–and the mean energy $\tilde{E}_{\mathrm{avg}}\left(m\right)$. And this mean energy is a graph-independent quantity, given by Eq. 18 (as modified by our filtering process, of course). Since MAX-3-XORSAT is an odd cost function, and our filtering choices are sign preserving, the density of states will be approximately Gaussian distributed around a peak at $E=0$ and $m=N/2$ (with width $O\left(\sqrt{N}\right)$), and one can argue from the law of large numbers that significant fluctuations about this average are proportionally rarer (and thus less relevant) with increasing $m$. Further, for random $H_{P}$ drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$, the largest possible string-to-string deviations from the average Eq. 18, uncorrelated with $G$, will scale as $O\left(N/\sqrt{d_{C}}\right)$ and are thus progressively more suppressed the more approximation-hard our random instances become. Distant configurations at each $m$ whose energy exhibits extensive deviations from Eq. 18 are thus a superpolynomially small fraction of the total, with global magnitude suppressed by a factor proportional to $1/\sqrt{d_{C}}$. Coupled with the regularization of the low energy landscape (which can influence the relative weight of configurations that have particularly high or low intermediate energies that must be crossed to reach them, starting from $G$), we therefore assert:
Claim 2 (accuracy of mean field calculations): for a random $H_{P}$ drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$, with high probability, but not guaranteed for all instances, the minimum gap $\Omega_{0}$ in the first stage of filtered optimization with $x\geq 2$ and $w\leq 1$ is well approximated by $\Omega_{\mathrm{mf}}\left(N\right)$, the result computed using a mean field cost function (Eq. 18) for $H_{P}$.
The ability to predict $\Omega_{0}$ from mean field analysis is a powerful tool, since we can massively pare down the Hilbert space to just $N+1$ states by employing permutation symmetry Jörg et al. (2010), and predict $\Omega_{\mathrm{mf}}\left(N\right)$ using exact diagonalization for thousands of spins without much trouble. This allows us to reliably access the real asymptotic behavior of this system, at least in the regime for which we claim these approximations are valid. And we note that, given the arguments above, these approximations become progressively better the further we are into the approximation-hard regime. The resulting functions, plotted in FIG. 5, are well behaved, and it is in their structure that the need for the warping transformation becomes obvious. Namely, for $w=1$ and any value of $x\geq 1$, the transition into the quantum spin glass ground state is first order and $\Omega_{\mathrm{mf}}\left(N\right)$ decays exponentially in $N$. For $x=2$ and $w=1$ the derived scaling exponent is smaller than for the unfiltered problem with $x=w=1$, but the decay of $\Omega_{\mathrm{mf}}\left(N\right)$ is still obviously exponential.
The crossover to polynomial scaling occurs when we choose a warping exponent $w<1$, or more specifically, $w<2/p$ for a $p$-body spin glass cost function. This is inspired by the observation that, for the sign preserving fractional power implemented in Eq. 9, the first order transition becomes second order when $w<2/p$ (at exactly $2/p$ the scaling may be stretched exponential for our sign-symmetric cost function), effectively replacing $p$ with $wp$ in Eq. 18. The intuitive mechanism for the removal of exponential decay is that higher energy states are pushed closer to $G$ in energy, in a nontrivial way that does not alter the overall normalization of a difference of $-N$ between $G$ and random states. This in turn increases their weight in $\left|G_{D}\right>$; we think it notable that the largest fractional changes to the magnitudes of the eigenvalues of $H_{\mathrm{cost}}^{x,w}$ are to energies near zero, where of course the density of states is exponentially largest. Note that this sequence of transformations (with $A=1$ in Eq. 8) does not change the energetic hierarchy of any classical states, only their scales; local and global minima of $H_{P}$ are still local and global minima of $H_{\mathrm{cost}}^{x,w}$.
Of course, if we choose $w$ to be too small, we expect performance to worsen, as when $w\to 0$ we lose the ability to discriminate $G$ from states with even tiny negative (but nonzero) energies. In real disordered problems this could manifest, among other issues, as worsening TFC, as local minima are pushed closer to $G$ and it is therefore easier (in principle) for transverse field corrections to push them below the ground state. As remarked earlier, and made increasingly clear over the course of this section, we see TFC as a much more fundamentally difficult problem than first order transitions, but not one that a priori limits quantum advantage for approximation given worst case classical thresholds that approach random guessing.
Choosing $w$ to be too small can also increase incidental performance degradations, as defined in section II.5, such as local heating, since the local excitation energies about $G$ continuously decrease with decreasing $w$. We therefore want to choose $w$ to be small enough to clearly be out of the exponential decay regime for $\Omega_{\mathrm{mf}}\left(N\right)$, but not smaller; empirically, values in the range of 0.5-0.65 work well for this purpose, depending on protocol choice. For the simple mean-field $p$-spin ferromagnet considered in Jörg et al. (2010), the lower limit to $w$, below which further reductions are counterproductive, can be established as $\simeq 1/p$. The local gap will steadily decrease at this point (leading to local heating unless runtime increases) without any corresponding improvement in the global minimum gap. For the disordered problems we consider here, we expect it to be somewhat higher than this, with an optimal value likely controlled by the clustering energy, though we do not presently have the mathematical tools to set an approximate bound. For concreteness, we choose $w=0.6$ in this work, but other values below $2/3$ are effective as well. For $w=0.6$, as seen in FIG. 5, $\Omega_{\mathrm{mf}}\left(N\right)\propto N^{-0.23}$ in numerical fitting. The entropic barrier is permeable.
This concludes the first stage of our algorithm, the stage which required the most conceptual work and analysis and which introduced a novel speedup mechanism, at least in our estimation. Having found the ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$, we now want to convert that into a good approximation, or even solution, to $H_{P}$, our underlying problem Hamiltonian, and to do so, we turn to stages 2 and 3.
Figure 5: Mean field minimum gap $\Omega_{\mathrm{mf}}\left(N\right)$ in the first stage of MSFO, for nonlinear fold parameter $x=2$, transverse field strength $\kappa=1$, and warp parameter $w=\left\{1,0.83,0.75,0.7,2/3,0.6\right\}$ (blue through brown, and monotonically increasing). For $w>2/3$, $\Omega_{\mathrm{mf}}\left(N\right)$ displays clearly exponential decay with $N$, with an exponent that steadily decreases with decreasing $w$. Conversely, for $w=0.6$ the best fit is to simple polynomial decay, $\Omega_{\mathrm{mf}}\left(N\right)\propto N^{-0.23}$, showing that the spectral warping transformation is capable of softening the exponential decay of the paramagnet-spin glass transition in this problem to low-order polynomial scaling.
IV.2.2 Second and third stages
To find solutions to $H_{P}$, we need to ramp down our transverse field $H_{D}$ to zero. The two filtering transformations we nested to minimize instance-dependent properties and eliminate the exponentially scaling first order transition both have the effect of reducing the local gap around $G$, and if we want to avoid local heating as we ramp the field down, we need to ramp the field down over a time $T_{2}\left(N\right)$ growing faster than $O\left(N^{2}\right)$ as the average local gap scales as $1/N$, with its minimum value across the lattice potentially decaying more quickly given variations in the density of constraints. While this doesn’t of course erase an exponential speedup, if getting into the basin of attraction of $G$, or finding any states sufficiently close to it in energy, takes exponential time classically, it’s suboptimal and can easily be improved upon. The room for improvement lies in the fact that despite $H_{\mathrm{cost}}^{x,w}$ being classically gapless, $H_{\mathrm{cost}}^{x,w}+\kappa H_{D}$ has an $O\left(1\right)$ gap on either side of the phase transition, as does the unfiltered $H_{P}+\kappa H_{D}$.
To achieve a faster time to solution, we therefore add two additional stages to the algorithm, though in principle they may be able to be combined into a single extra stage. The key observation motivating these stages is that $G$ is a gapped, presumably unique ground state of $H_{P}$, and the ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$ is spread across a much larger range of configurations, centered around $G$ and with the largest amplitudes in its basin of attraction. If however we are well below $\kappa_{c}$, which we expect to be for $\kappa=1$ and $p=3$, then if we smoothly reduce $x$ to 1 and increase $w$ to 1, turning off the filtering transformations, then there is no a priori reason that a phase transition should be crossed, and indeed, the gap does not close at the mean field level when this is done. Assuming this is done sufficiently slowly–and here, we use the same scaling with $N$ as the first stage, though this is likely more than necessary as the gap doesn’t close–we are now in the quantum spin glass ground state of $H_{P}+\kappa H_{D}$. Had we not implemented any of the filtering transformations, and instead attempted to find the ground state of $H_{P}$ directly, this step would have taken exponential time, as predicted in Eq. 25 and confirmed later in our TAQC simulations.
Finally, in stage 3, we ramp $\kappa$ down to zero and measure the system. For large $d_{C}$, the classical problem is locally gapped with high probability, since the variance of the local gap decreases with increasing $d_{C}$, and so this step too can be done quickly. Assuming we remained in the ground state in stages 1 and 2, we will reach the ground state in this stage and can halt and measure the system. That said, while at large $d_{C}$ and small $\epsilon$ the classical $G$ is locally gapped with high probability, this is obviously not the case for all possible instances, and instance-dependent variations can impede performance in these steps. We also expect that, in practice, stages 2 and 3 can be combined into a single step (where $\kappa\to 0$, $w\to 1$ and $x\to 1$ all occur simultaneously), but focus on the full three stage version in this work.
IV.2.3 Time to solution scaling
Assuming claims 1 and 2 hold, it is straightforward to predict the time to solution scaling of MSFO. Specifically, assuming that the first stage is the most computationally challenging step, $P\left(t_{F}\right)$ is constant if we choose
$t_{F}\left(N\right)\propto W\left(N\right)\Omega_{\mathrm{mf}}\left(N\right)^{-2}.$ (26)
For algorithm parameters $x=2$ and $w=0.6$, the asymptotic scaling of $\Omega_{\mathrm{mf}}\left(N\right)$ is $\Omega_{\mathrm{mf}}\left(N\right)\propto N^{-0.23}$, and assuming $W\left(N\right)\propto N$ this leads to an asymptotic runtime scaling of $O\left(N^{1.46}\right)$, again for the normalization that the ground state of $H_{\mathrm{cost}}$ is $-N$ and random states have mean energy zero. Given the polynomial scaling and transition stability assumption stated above, this could likely be further improved through schedule fine tuning, or through choosing a slightly smaller value of $w$. We note that this is an asymptotic estimate, and for the relatively small $N$ values accessible to numerical simulation, the scaling remains polynomial with roughly the same degree, but there are additional complexities to consider to obtain accurate predictions. We take up this issue in section VI.3.
To get the total gate count of the algorithm, we need to incorporate a few additional factors. First is a factor of $O\left(N_{C}\right)$ that comes from the need to guess the satisfied fraction $\epsilon$ to set the energy scale for the two nonlinear folding and warping transformations; there are of course only $O\left(N_{C}\right)$ possible values that this can take so at worst case we can just rerun the entire algorithm $O\left(N_{C}\right)$ times with different choices and take the best result. There is likewise a factor of $O\left(N_{C}\times{\rm polylog}\left(N\right)\right)$ that comes from computing $H_{\mathrm{cost}}=F_{w}\left(F_{f}\left(H_{P},N,1,x\right),N,w\right)$ at each timestep using ancillary qubits, as sketched in FIG. 2. Finally, to avoid heating from Trotter error, we may asymptotically need to reduce the timestep $dt$ with increasing $N$, or make the circuit more complex at each timestep. In either case we expect a worst case factor sublinear in $N$. Putting all these considerations together, we arrive at the time to solution scaling of $O\left(N^{2}N_{C}^{2}\times{\rm polylog}\left(N\right)\right)$ in our core conjecture.
IV.3 Spectrally folded trial minimum annealing performance analysis
Our multi-stage algorithm employed a sequence of composed filters, with the goal of reaching a single ground state (or at least, states near it in energy and/or hamming distance). If we instead focus only on finding states below a given approximation threshold $A$, the warping transformation is not necessary and a simple linear folding procedure is sufficient to find approximate solutions quickly. To further support our core conjecture, we now present a second algorithm, which employs only a linear folding transformation and, starting from a poor quality local minimum found through simple classical methods, can find good approximate solutions to $H_{P}$ through collective quantum tunneling processes. This is the spectrally folded trial minimum annealing algorithm (TMA for the remainder of this section) presented in section III.4. The performance of these two algorithms is comparable, but they arrive at solutions through completely different mechanisms and the calculations predicting scaling are substantially different.
As outlined in section III.4, the TMA calculation starts by employing a simple classical algorithm to find a local minimum state of $H_{P}$. In this work, we use the quasi-greedy algorithm in section V.2 for this purpose, but essentially any simple method can be used; all that matters is that it’s a local minimum and the initial energy of this state does not matter. We assume that our problem $H_{P}$ is approximation-hard in practice (as if it isn’t, there’s no reason to use any of these new methods), and the energy of our initial state is much higher than that our approximation target $-AN$, for control parameter $A\leq 1$. Once again we assume we have normalized $H_{P}$ so that its ground state energy is $-N$. We assume that it is uncorrelated with $G$ and thus $M\sim N/2$ flips away from it. We let this initial state be $\left|L\right>$, and apply a classical lowering Hamiltonian $H_{L}$ to lower it into competition with the folded ground state band. We let the set of all states in or near this band be $\mathcal{T}$. The transverse field is then turned on, $H_{L}$ is slowly ramped down to zero, and in the process, states in the folded ground state band are found through collective tunneling processes.
Our performance prediction for TMA is set up as a stability calculation: specifically, we ask if the dressed state $\left|L_{D}\right>$ ($\left|L\right>$ dressed by transverse field corrections) is asymptotically stable as it crosses the band center of $\mathcal{T}$. If is unstable and decays, as the only states in energetic competition are those in $\mathcal{T}$, that means the algorithm has found an approximate solution to $H_{P}$. The stability calculation itself uses a standard Fermi’s Golden Rule analysis and predicts the decay rate by multiplying the average squared per-state decay matrix element $\Omega_{0,Lj}^{2}\right>_{j}$ by the total number of states in $\mathcal{T}$; if the exponential decay of $\Omega_{0,Lj}^{2}\right>_{j}$ is balanced or exceeded by the exponential growth of targets in $\mathcal{T}$, $\left|L_{D}\right>$ is unstable and solutions will be found quickly. Given that Eq. 18 is asymptotically exact, predicting the density of states in $\mathcal{T}$ is easy, [^8] Throughout this calculation, we assume a single ground state $G$. The way the calculation is structured, additional deep minima simply add additional targets and make the problem easier; our predictions here are thus in some sense a worst case estimate, at least for random $H_{P}$ drawn from $\mathcal{P}\left(N,N_{C},\epsilon\right)$ as specified in our core conjecture. but predicting the average per-state tunneling rate required the development of a new high-order perturbation theory framework. We begin that calculation now.
IV.3.1 Calculation of the per-state tunneling rate
As in the MSFO calculation, the key ingredients of the per-state tunneling rate are the average cost per flip away from either minimum and the transverse field strength $\kappa$. Given a ground state energy $-N$, we assume that $\kappa<\kappa_{c}$, the critical value for which $\left|L_{D}\right>$ would transition to a paramagnetic ground state, estimated to be $\kappa_{c}\simeq 1.29$ earlier. We need to choose $\kappa$ to be less than this value to ensure we remain in the dressed problem phase (DPP) and the calculations performed here are valid.
The average cost per flip curve takes a little more work to determine, but is ultimately also straightforward. As our goal in this section is to compute the average multiqubit tunneling matrix element between $\left|L_{D}\right>$ and a randomly chosen state $\left|j_{D}\right>$ near the fold in $\mathcal{T}$, starting from one of the two states we need to differentiate between flips that move closer to the other state, and flips that move further away. We call these processes flips of primary and secondary spins, respectively; a primary spin is one whose value changes between the classical strings $\left|L\right>$ and $\left|j\right>$, and a secondary spin is one that does not. We note that the typical state in $\mathcal{T}$ is $M\simeq N/2$ flips away from $\left|L\right>$ (we address the issue of varying Hamming distance later on), and assume that $\left|L_{D}\right>$ crosses the band center of $\mathcal{T}$, where the density of states is exponentially largest, when the classical energy of both states is $\simeq-N$. We revisit this assumption later on.
The mean cost per flip curve away from $\left|L\right>$ is trivially given by Eq. 18, as $H_{L}$ is a random 3-XORSAT problem Hamiltonian. For flips away from $\left|T\right>$, we note that the average state at or near the fold is $x_{A}$ flips away from $G$. To find $x_{A}$ we can invert Eq. 18 to find the mean number of flips $x_{A}N$ for which $E\left(x_{A}N\right)\right>=-AN$. To be specific,
$x_{A}=\frac{1-A^{1/3}}{2}.$ (27)
We can therefore assume that the typical state in $\mathcal{T}$ is $x_{A}N$ flips away from $\left|G\right>$. If we consider the sequences of primary spin flips connecting $\left|j_{D}\right>$ and $\left|L_{D}\right>$, the typical flip sequence starts $x_{A}N$ flips away from $\left|G\right>$, and notice that with probability $1-x_{A}$ an additional random flip towards $\left|L\right>$ will also move closer to $\left|G\right>$. Taking all these effects and the division by $A$ into account, so that the bare unperturbed energy of both $\left|j\right>$ and $\left|L\right>$ when they cross are both $\sim-N$, a bit of algebra shows that for $y$ flips away from a ground state of the folded Hamiltonian, not only is the total average cost $\Delta E\left(y\right)$ $A$-independent, it is precisely equal to the cost given by Eq. (18). We note that all of the arguments about the statistical regularity of random-hypergraph problems from the previous section apply here as well.
We can are now ready to define the bare classical energy, with no corrections from transverse fields. We start from one of the two states, and consider a state which is $m+n$ random flips away from it. We let $m$ of these flips be ones which move toward the other minimum (e.g. reduce the Hamming distance to it), and the $n$ flips be flips that move away from it. Then the bare average energy $E_{m,n}^{(0)}$ is given by:
@@ IV.3.1 Calculation of the per-state tunneling rate
  • We divide the system between primary spins, which flip between the classical minima, and secondary spins, which do not. We then compute the dressed states $\left|L_{D}\right>$ and $\left|j_{D}\right>$ that comprise all the primary flip sequences up to order $M/2$ away from each minimum. It is at this order that the two states have nonzero overlap.
@@
$\left|L_{D}\right>\equiv\left|L\right>+\sum_{j}\frac{\kappa}{u_{1,0}}X_{j}\left|L\right>+\sum_{j,k\left(j\neq k\right)}\frac{2\kappa^{2}}{u_{1,0}u_{2,0}}X_{j}X_{k}\left|L\right>$
$+\sum_{j,k,l\left(j\neq k\neq l\right)}\frac{3!\kappa^{3}}{u_{1,0}u_{2,0}u_{3,0}}X_{j}X_{k}X_{l}\left|L\right>+...$ (30)
If we ignore the local band structure, the expression for $\left|j_{D}\right>$ is functionally identical. Neglecting this structure is expected to underestimate the total collective tunneling rate; we revisit this issue in section C. We stop our expansion at order $M/2$, which is the lowest nontrivial order needed to connect the states. For simplicity we assume $M$ is even though the argument is easy to generalize to odd $M$ as well. Note that this state is not normalized, and in fact the norm of the state written above is exponentially large, so we will need to incorporate normalization corrections into the definition of the states. Thanks to the dressing of the states, we obtain a primary-spin energy splitting
$\frac{1}{2}\left(\left<L_{D}\right|+\left<j_{D}\right|\right)H_{P}\left(\left|L_{D}\right>+\left|j_{D}\right>\right)=2\Omega_{0}^{(p)},$
@@
$\mathcal{N}^{(p)}$ $=$ $\left(1+\sum_{k=1}^{M/2}\binom{M}{k}\left(\kappa^{k}k!\prod_{l=1}^{k}\frac{1}{u_{l,0}}\right)^{2}\right)^{-1}.$ (32)
@@
We now turn to the secondary spins. To introduce secondary spin corrections, consider a single secondary spin $j$ out of the $\left(N-M\right)\sim N/2$ total, whose bit value is the same in both classical minima. Since the same transverse field is acting on it as all other spins, when we consider the sum of all processes that connect the two minima, we can now divide them between those where $M/2$ primary spins flip from each minima to meet in the middle, with secondary spin $j$ unchanged, and a new set of processes where spin $j$ flips starting from each minimum and the two wavefunctions overlap at the set of states where $M/2$ primary spins have flipped along with $j$. The first set of processes is what was considered in Eq. IV.3.1; the second is new, and we want to calculate its matrix element. It is most useful to express these matrix elements as a ratio of the new term to the original, primary-spin-only process, since both decay exponentially in $M$. Let the primary spin perturbative matrix element to reach $M/2$ flips be $\xi_{M/2}$, so that
@@
Taking all of these effects into account yields a highly accurate prediction of the minimum gap scaling for a wide range of values for $p$ and $\kappa$, with only an $O\left(1\right)$ discrepancy in the prefactor and few percent discrepancies in the scaling exponent (empirically, Eq. IV.3.1 tends to slightly overestimate the decay compared to the exponent extracted from numerical diagonalization). We refer to the Appendix A for more details.
@@
IV.3.2 Achievable approximation ratio with spectrally folded trial minimum annealing